# SemPLeS: Semantic Prompt Learning for Weakly-Supervised Semantic Segmentation

### v2

@@ Abstract
Weakly-Supervised Semantic Segmentation (WSSS) aims to train segmentation models using image data with only image-level supervision. Since precise pixel-level annotations are not accessible, existing methods typically focus on producing pseudo masks for training segmentation models by refining CAM-like heatmaps. However, the produced heatmaps may capture only the discriminative image regions of object categories or the associated co-occurring backgrounds. To address the issues, we propose a Semantic Prompt Learning for WSSS (SemPLeS) framework, which learns to effectively prompt the CLIP latent space to enhance the semantic alignment between the segmented regions and the target object categories. More specifically, we propose Contrastive Prompt Learning and Prompt-guided Semantic Refinement to learn the prompts that adequately describe and suppress the co-occurring backgrounds associated with each target object category. In this way, SemPLeS can perform better semantic alignment between object regions and the associated class labels, resulting in desired pseudo masks for training the segmentation model. The proposed SemPLeS framework achieves SOTA performance on the standard WSSS benchmarks, PASCAL VOC and MS COCO, and shows compatibility with other WSSS methods. The source codes are provided in the supplementary.
Keywords:
Weakly-Supervised Semantic Segmentation CLIP Prompt
@@ 1 Introduction
Semantic segmentation aims to classify every pixel in images to identify object categories and the associated regions, which can benefit various applications in the real world [52, 47, 73]. While promising results have been presented by fully-supervised approaches [38, 44, 5, 6, 75, 7, 76], collecting pixel-level annotations could be time-consuming and expensive, and therefore limits the scalability and practicality of fully-supervised methods. To address this issue, Weakly-Supervised Semantic Segmentation (WSSS) has emerged as an alternative approach to train segmentation models with only coarse or incomplete annotations such as bounding boxes [28], scribbles [39], points [3], or image-level labels. Among these annotation forms, image-level labels which indicate the presence or absence of certain object categories are commonly used due to the efficiency in data collection and the availability in various benchmark image datasets. Since precise annotations of object positions are not observed, learning to localize and segment object categories from image-level supervision is particularly challenging. Most existing methods [4, 61, 54, 67, 27, 31] focus on producing pseudo ground truth masks by learning CAM-like heatmaps [78, 56] with class labels as discriminative supervision. Despite the shown efficacy, the learned CAMs may still miss relevant regions of target object categories and fail to cover the entire object. Furthermore, co-occurring backgrounds associated with certain object categories may also be falsely activated (e.g., river or lake in an image of a car). Consequently, learning precise image regions that align with the semantics of target objects from weak supervision remains a challenging task.
@@
With the rapid growth in the amount of image and text data in recent years, several vision-language models [45, 11, 50] have been proposed to bridge the underlying semantics between the two modalities. Given that both the images and the associated class labels (object names) are available in the setting of WSSS, the underlying image-text semantics from the CLIP [50] latent space can be leveraged to enhance the quality of CAMs and pseudo masks. Recent CLIP-based methods [64, 41, 71, 48, 68] mainly focus on designing text prompts or prompt learning techniques for the text encoder. Despite the effectiveness demonstrated, they either consider only the foreground class prompts, or rely on general background prompts (e.g., “a photo of river”, “a photo of lake”, etc.) defined by additional manual efforts and heuristic human knowledge, as shown in Fig. 1 (a). Moreover, such manually-defined prompts may not fully exploit the knowledge in the CLIP latent space.
Figure 1: (a) Previous CLIP-guided WSSS methods require additional manual prompt engineering effort involved with heuristic human knowledge, while (b) our proposed SemPLeS framework automatically learns prompts embedded with class-associated semantic knowledge discovered from the CLIP latent space without any manual effort.
@@
In this paper, we aim to fully exploit the CLIP latent space to benefit the weakly-supervised semantic segmentation problem without manual prompting. To achieve this goal, we propose a Semantic Prompt Learning for WSSS (SemPLeS) framework to learn prompts embedded with class-associated semantic knowledge discovered from the CLIP latent space, as shown in Fig. 1 (b), where the learned prompts can enhance the semantic alignment between the segmented regions and the target object categories with image-level labels. More specifically, we perform image-text contrastive learning under the guidance of CLIP and train a mask generator to generate class activation maps. However, such produced masks might not be sufficiently precise, and the co-occurring backgrounds associated with the object categories may be falsely activated. To alleviate this problem, we uniquely present Contrastive Prompt Learning and Prompt-guided Semantic Refinement to suppress class-associated background regions. In Contrastive Prompt Learning, we learn prompts to capture co-occurring backgrounds from images and labels. Without manually defining the background texts, our learned prompts would properly describe the backgrounds associated with target object categories. Under the guidance of our class-associated background prompts, we further suppress co-occurring backgrounds from the activation maps via Prompt-guided Semantic Refinement. With the above designed learning strategy, the semantic matching between object regions and the associated class labels will be enhanced, resulting in desired activation maps for segmentation purposes. The proposed SemPLeS framework achieves SOTA performance on the standard WSSS benchmarks, PASCAL VOC and MS COCO. Moreover, our SemPLeS framework can integrate and consistently improve other WSSS methods including CNN-, Transformer-, and foundation model based ones, confirming its effectiveness and compatibility.
@@
  • We propose a novel Semantic Prompt Learning for WSSS (SemPLeS) framework, which fully exploits the CLIP latent space to benefit the weakly-supervised semantic segmentation without manual prompting. Moreover, our SemPLeS framework shows compatibility with other WSSS methods, including CNN-, Transformer-, and foundation model based ones.
  • In SemPLeS, we present Contrastive Prompt Learning to learn prompts embedded with class-associated semantic knowledge. With no need to manually define background texts, our learned prompts would represent co-occurring backgrounds associated with distinct object categories.
  • With the derived prompts, our Prompt-guided Semantic Refinement learns to suppress co-occurring backgrounds while enhancing the semantic alignment between object regions and the associated class labels, resulting in precise pseudo masks and state-of-the-art segmentation performance in WSSS.
@@ 2 Related Works
Weakly-Supervised Semantic Segmentation (WSSS):
Existing WSSS approaches typically follow a three-stage learning process. Firstly, the image-level labels are utilized as supervision to generate initial Class Activation Maps (CAMs) [78, 56]. Secondly, the initial CAMs are refined by using dense CRF [30] or pixel affinity-based methods [2, 1] to obtain pseudo masks. Lastly, the refined pseudo masks are further exploited to train a segmentation model. Among all the stages, producing precise CAMs (i.e., the first stage) is the main focus of WSSS, and various approaches have been proposed to improve the quality of CAMs [21, 9, 14, 65, 19, 26, 34, 32, 72, 63, 35]. With the rapid development and the success of vision transformers [18], recent approaches [54, 67, 53, 55, 16, 68, 49, 82] generate finer activation maps based on the patch-level affinity learned from the attention layers. Very recently, several works [10, 25, 59, 13, 71] exploit foundation segmentation models (e.g., SAM [29]) to enhance the quality of the pseudo masks. On the other hand, there are also end-to-end WSSS works [55, 62] which do not require multiple training stages, yet their performances are inferior to standard 3-stage methods. In general, most WSSS methods learn CAMs through object classification, overlooking the textual semantics of class labels. Instead, our method exploits vision-language models to discover class-associated semantic knowledge, therefore producing high-quality CAMs for segmentation.
CLIP-based Semantic Segmentation:
@@ 3 Proposed Method
Figure 2: An overview of our proposed SemPLeS framework. We first introduce (a) Segment-Label Matching, which leverages image-text contrastive learning to produce initial object masks $M$ from the mask generator $S$. To achieve class-associated mask refinement and produce the refined mask $M^{\prime}$, we propose (b) Contrastive Prompt Learning to automatically learn prompts $p_{k}$ embedded with semantic knowledge, followed by (c) Prompt-guided Semantic Refinement to suppress co-occurring backgrounds associated with each category $k$.
@@ 3.1 Problem Formulation and Model Overview
3.2 Semantic Prompt Learning for WSSS (SemPLeS)
Segment-Label Matching:
@@ 3.2 Semantic Prompt Learning for WSSS (SemPLeS)
Given an input image $X$, our mask generator $S$ is designed to produce soft foreground masks $M=S(X)$ for target object categories. Since pixel-wise annotations are not available, we choose to leverage vision-language models to guide the learning of our mask generators from image-level supervision. To be more precise, we exploit the joint latent space for images and texts from CLIP to match the object regions and the associated text labels. To achieve this, an image-text triplet (i.e., foreground-background-text) would be formulated to perform contrastive learning, as illustrated in Fig. 2 (a). For the $k$th ground truth category which presents in the input image $X$ (i.e., $y_{k}=1$), we derive the foreground image $X^{f}_{k}=M_{k}\cdot X$ by applying the $k$th predicted mask $M_{k}$ to the original image $X$. Similarly, we reverse the predicted mask to obtain the background regions $X^{b}_{k}=(1-M_{k})\cdot X$. As for the text input $t_{k}$, we adopt the common prompt template “a photo of {}” filled with the $k$th class name to describe the category of interest. With the triplet [$X^{f}_{k}$, $X^{b}_{k}$, $t_{k}$] serving as the input of the image encoder $E_{I}$ and text encoder $E_{T}$ pre-trained by CLIP, we perform image-text contrastive learning to maximize the cosine similarity between $X^{f}_{k}$ and $t_{k}$ for the foreground, while the similarity of $X^{b}_{k}$ and $t_{k}$ would be minimized to repel the background. Therefore, our matching loss $L_{match}$ would be formulated as follows:
$L_{match}=\mathbb{E}_{X}-\left[log(sim(v^{f}_{k},u^{f}_{k}))+\lambda_{b}\cdot log(1-sim(v^{b}_{k},u^{f}_{k}))\right],$ (1)
$\text{where}\quad v^{f}_{k}=E_{I}(X^{f}_{k}),\quad v^{b}_{k}=E_{I}(X^{b}_{k})\quad\text{and}\quad u^{f}_{k}=E_{T}(t_{k}).$
@@
Here, $\lambda_{b}$ is the loss weight for repelling backgrounds and $sim$ refers to cosine similarity. Note that we keep the image encoder $E_{I}$ and the text encoder $E_{T}$ frozen during training and preserve the latent space learned from CLIP to avoid potential overfitting. With the above Segment-Label Matching, our mask generator $S$ is encouraged to segment object regions that align with the associated text labels. However, as noted above, such masks learned from image-level supervision are still coarse, and may falsely include co-occurring backgrounds associated with certain object categories. Therefore, the above image-text matching is not sufficient for segmentation.
Contrastive Prompt Learning:
@@
To address the coarse mask issues, the previous language-guided approach [64] exploits vision-language models to refine the masks with manual prompting techniques. However, these methods require additional prompt engineering efforts with human knowledge involved. Moreover, manual prompting may not be able to fully exploit vision-language representation space. To tackle these problems, we propose Contrastive Prompt Learning (Fig. 2 (b)) to learn prompts embedded with semantic knowledge from vision-language models, facilitating the following object mask refinement. Different from the previous work, we employ a sequence of learnable prompts $p_{k}$ as the input of the text encoder $E_{T}$ to describe backgrounds for each distinct category $k$. Specifically, to align the prompts $p_{k}$ with the background image $X^{b}_{k}$, we maximize the similarity of their representations in the latent space by proposing $L^{I}_{prompt}$. On the other hand, to avoid describing the foreground objects, we encourage the similarity between $u^{b}_{k}$ and $u^{f}_{k}$ to be low with the proposed $L^{T}_{prompt}$. Thus, we propose $L_{prompt}$ as below:
@@
$=\mathbb{E}_{X}\left[-log(sim(u^{b}_{k},v^{b}_{k}))+\lambda_{T}\cdot log(sim(u^{b}_{k},u^{f}_{k}))\right],$
$\text{where}\quad u^{b}_{k}=E_{T}(p_{k}),\quad v^{b}_{k}=E_{I}(X^{b}_{k})\quad\text{and}\quad u^{f}_{k}=E_{T}(t_{k}).$
Here, the mask generator $S$ is fixed and $p_{k}$ is the only trainable part for loss $L_{prompt}$, and $\lambda_{T}$ is the loss weight for minimizing the similarities to the object categories. Once the above learning is complete, our prompts $p_{k}$ would represent co-occurring backgrounds for each category $k$ without requiring manually-defined background prompts, and is therefore preferable to existing CLIP-based methods [64, 41, 71, 48, 68]. In addition, our Contrastive Prompt Learning aims to capture class-associated backgrounds for segmentation purposes, rather than replacing general text templates like “a photo of {}” for classification tasks as previous prompt learning methods [81, 80, 24] do.
Prompt-guided Semantic Refinement:
Finally, to suppress co-occurring background regions from the object mask $M$, our SemPLeS framework exploits the previously derived background prompts $p_{k}$ to perform Prompt-guided Semantic Refinement (Fig. 2 (c)). More specifically, we encourage our mask generator $S$ to produce refined masks $M^{\prime}$ by excluding the semantic knowledge embedded in the background prompts $p_{k}$, while the objectives introduced in Eq. 1 are retained to match the class labels. Hence, the refinement loss $L_{refine}$ and the total loss function $L_{total}$ are defined as follows:

### v3

# Semantic Prompt Learning for Weakly-Supervised Semantic Segmentation

@@ Abstract
Weakly-Supervised Semantic Segmentation (WSSS) aims to train segmentation models using image data with only image-level supervision. Since precise pixel-level annotations are not accessible, existing methods typically focus on producing pseudo masks for training segmentation models by refining CAM-like heatmaps. However, the produced heatmaps may capture only the discriminative image regions of object categories or the associated co-occurring backgrounds. To address the issues, we propose a Semantic Prompt Learning for WSSS (SemPLeS) framework, which learns to effectively prompt the CLIP latent space to enhance the semantic alignment between the segmented regions and the target object categories. More specifically, we propose Contrastive Prompt Learning and Prompt-guided Semantic Refinement to learn the prompts that adequately describe and suppress the co-occurring backgrounds associated with each object category. In this way, SemPLeS can perform better semantic alignment between object regions and class labels, resulting in desired pseudo masks for training segmentation models. The proposed SemPLeS framework achieves competitive performance on standard WSSS benchmarks, PASCAL VOC 2012 and MS COCO 2014, and shows compatibility with other WSSS methods.
**footnotetext:  Work done during an internship at NVIDIA.
@@ 1 Introduction
Semantic segmentation aims to classify every pixel in images to identify object categories and the associated regions, which can benefit various applications in the real world [53, 48, 75]. While promising results have been presented by fully-supervised approaches [39, 45, 6, 7, 77, 8, 78], collecting pixel-level annotations could be time-consuming and expensive, and therefore limits the scalability and practicality of fully-supervised methods. To address this issue, Weakly-Supervised Semantic Segmentation (WSSS) has emerged as an alternative approach to train segmentation models with only coarse or incomplete annotations such as bounding boxes [29], scribbles [40], points [3], or image-level labels. Among these annotation forms, image-level labels which indicate the presence or absence of certain object categories are commonly used due to the efficiency in data collection and the availability in various benchmark image datasets. Since precise annotations of object positions are not observed, learning to localize and segment object categories from image-level supervision is particularly challenging. Most existing methods [4, 61, 55, 68, 28, 32] focus on producing pseudo ground truth masks by learning CAM-like heatmaps [80, 57] with class labels as discriminative supervision. Despite the shown efficacy, the learned CAMs may still miss relevant regions of target object categories and fail to cover the entire object. Furthermore, co-occurring backgrounds associated with certain object categories may also be falsely activated (e.g., rails in a photo of a train). Consequently, learning precise image regions that align with the semantics of target objects from weak supervision remains a challenging task.
@@
With the rapid growth in the amount of image and text data in recent years, several vision-language models [46, 11, 51] have been proposed to bridge the underlying semantics between the two modalities. Given that both the images and the associated class labels (object names) are available in the setting of WSSS, the underlying image-text semantics from the CLIP [51] latent space can be leveraged to enhance the quality of CAMs and pseudo masks. Recent CLIP-based methods [65, 42, 72, 49, 69] mainly focus on designing text prompts or prompt learning techniques for the text encoder. Despite the effectiveness demonstrated, they either consider only the foreground class prompts, or rely on general background prompts (e.g., “a photo of rail”, “a photo of road”, etc.) defined by additional manual efforts and heuristic human knowledge, as shown in Fig. LABEL:figure:teaser (a). Moreover, such manually-defined prompts may not fully exploit the knowledge in the CLIP latent space.
@@
In this paper, we aim to fully exploit the CLIP latent space to benefit the weakly-supervised semantic segmentation problem without manual prompting. To achieve this goal, we propose a Semantic Prompt Learning for WSSS (SemPLeS) framework to learn prompts embedded with class-associated semantic knowledge discovered from the CLIP latent space, as shown in Fig. LABEL:figure:teaser (b), where the learned prompts can enhance the semantic alignment between the segmented regions and the target object categories with image-level labels. More specifically, we perform image-text contrastive learning under the guidance of CLIP and train a mask generator to generate class activation maps. Such produced object masks, however, might not be sufficiently precise, and the co-occurring backgrounds associated with the object categories may be falsely activated. To alleviate this problem, we uniquely present Contrastive Prompt Learning and Prompt-guided Semantic Refinement to suppress class-associated background regions. In Contrastive Prompt Learning, we learn prompts to capture co-occurring backgrounds from images and class labels. Without manually defining the background texts, our learned prompts would properly describe the backgrounds associated with each object category. Under the guidance of our learned class-associated background prompts, we further suppress co-occurring backgrounds from the activation maps via Prompt-guided Semantic Refinement. With the above-designed learning strategy, the semantic matching between object regions and the associated class labels will be enhanced, resulting in precise pseudo masks desired for training segmentation networks. The proposed SemPLeS framework achieves competitive performance on standard WSSS benchmarks, PASCAL VOC 2012 and MS COCO 2014. Moreover, our proposed SemPLeS framework can integrate and improve other WSSS methods including CNN-, Transformer-, and foundation model-based ones, confirming its effectiveness and compatibility.
@@
  • We propose a novel Semantic Prompt Learning for WSSS (SemPLeS) framework, which fully exploits the CLIP latent space to benefit the weakly-supervised semantic segmentation without manual prompting. Additionally, our SemPLeS framework shows compatibility with other WSSS methods, including CNN-, Transformer-, and foundation model-based ones.
  • In SemPLeS, we present Contrastive Prompt Learning to learn prompts embedded with class-associated semantic knowledge. With no need to manually define background texts, our learned prompts would properly capture co-occurring backgrounds associated with distinct object categories.
  • With the derived prompts, our Prompt-guided Semantic Refinement learns to suppress co-occurring backgrounds while enhancing the semantic alignment between object regions and the associated class labels, resulting in precise pseudo masks and competitive segmentation performance in WSSS.
@@ 2 Related Works
2.1 Weakly-Supervised Semantic Segmentation
Existing WSSS approaches typically follow a three-stage learning process. Firstly, the image-level labels are utilized as supervision to generate Class Activation Maps (CAMs) [80, 57]. Secondly, the CAMs are refined by using dense CRF [31] or pixel affinity-based methods [2, 1] to obtain pseudo masks. Lastly, the pseudo masks are further exploited to train segmentation networks. Among all the stages, producing precise CAMs (i.e., the first stage) is the main focus of WSSS, and various approaches have been proposed to improve the quality of CAMs [21, 9, 14, 66, 19, 27, 35, 33, 73, 63, 36]. With the rapid development and the success of vision transformers [18], recent approaches [55, 68, 54, 56, 16, 69, 50, 84] generate finer activation maps based on the patch-level affinity learned from the attention layers. Very recently, several works [10, 26, 59, 13, 72] exploit foundation segmentation models (e.g., SAM [30]) to enhance the quality of the pseudo masks. On the other hand, there are also end-to-end WSSS works [56, 62] which do not require multiple training stages, yet their performances are inferior to standard 3-stage methods. In general, most WSSS methods learn CAMs through object classification, overlooking the textual semantics of class labels. Instead, our method exploits vision-language models to discover class-associated semantic knowledge, therefore producing high-quality CAMs for segmentation.
2.2 CLIP-based Semantic Segmentation
@@ 3 Proposed Method
Figure 1: An overview of our proposed SemPLeS framework. We first introduce (a) Segment-Label Matching, which leverages image-text contrastive learning to train the mask generator $S$ and produce initial object masks $M$. Such derived masks are still coarse and may falsely include co-occurring backgrounds. To achieve class-associated mask refinement and produce the refined mask $M^{\prime}$, we propose (b) Contrastive Prompt Learning to automatically learn prompts $p_{k}$ embedded with semantic knowledge from the CLIP latent space, followed by (c) Prompt-guided Semantic Refinement to suppress co-occurring backgrounds associated with each category $k$.
@@ 3.1 Problem Formulation and Model Overview
3.2 Semantic Prompt Learning for WSSS
3.2.1 Segment-Label Matching
@@
Given an input image $X$, our mask generator $S$ is designed to produce foreground masks $M=S(X)$ for target object categories. Since pixel-wise annotations are not available, we choose to leverage vision-language models to guide the learning of our mask generators from image-level supervision. To be more precise, we exploit the joint latent space for images and texts from CLIP to learn object regions aligned with the associated text labels. To achieve this, an image-text triplet (i.e., foreground-background-text) would be formulated to perform contrastive learning, as illustrated in Fig. 1 (a). For the $k$th ground truth category which presents in the input image $X$ (i.e., $y_{k}=1$), we derive the foreground image $X^{f}_{k}=M_{k}\cdot X$ by applying the $k$th predicted mask $M_{k}$ to the original image $X$. Similarly, we reverse the predicted mask to obtain the background regions $X^{b}_{k}=(1-M_{k})\cdot X$. As for the text input $t_{k}$, we adopt the common prompt template “a photo of {}” filled with the $k$th class name to describe the category of interest. With the triplet [$X^{f}_{k}$, $X^{b}_{k}$, $t_{k}$] serving as the input of the image encoder $E_{I}$ and text encoder $E_{T}$ pre-trained by CLIP, we perform image-text contrastive learning to maximize the cosine similarity between $X^{f}_{k}$ and $t_{k}$ for the foreground, while the similarity of $X^{b}_{k}$ and $t_{k}$ would be minimized to repel the background. Therefore, our matching loss $L_{match}$ would be formulated as follows:
$L_{match}=$ $\mathbb{E}_{X}\left[-log(sim(v^{f}_{k},u^{f}_{k}))\right]+$ (1)
$\mathbb{E}_{X}\left[-\lambda_{b}\cdot log(1-sim(v^{b}_{k},u^{f}_{k}))\right],$
$\textit{where}\,\,v^{f}_{k}=E_{I}$ $(X^{f}_{k}),\,\,v^{b}_{k}=E_{I}(X^{b}_{k}),\,\,u^{f}_{k}=E_{T}(t_{k}).$
@@
Here, $\lambda_{b}$ is the loss weight for repelling backgrounds and $sim$ refers to cosine similarity. Note that we keep the image encoder $E_{I}$ and the text encoder $E_{T}$ frozen during training and preserve the latent space learned from CLIP to avoid potential overfitting. With the above Segment-Label Matching, our mask generator $S$ is encouraged to distinguish foregrounds and backgrounds with the associated text labels. However, as noted above, such masks learned from image-level supervision are still coarse, and may falsely include co-occurring backgrounds associated with certain object categories. Therefore, the above image-text matching is not sufficient to achieve precise segmentation.
3.2.2 Contrastive Prompt Learning
@@
To address the coarse mask issues, the previous language-guided approach [65] exploits vision-language models to refine the masks with manual prompting techniques. However, these methods require additional prompt engineering efforts with human knowledge involved. Moreover, manual prompting may not be able to fully exploit vision-language representation space. To tackle these problems, we propose Contrastive Prompt Learning (Fig. 1 (b)) to learn prompts embedded with semantic knowledge from vision-language models, facilitating the following object mask refinement. Different from the previous work, we employ a sequence of learnable prompts $p_{k}$ as the input of the text encoder $E_{T}$ to describe backgrounds for each distinct category $k$. Specifically, to align the prompts $p_{k}$ with the background image $X^{b}_{k}$, we maximize the similarity of their representations in the latent space via $L^{I}_{prompt}$. On the other hand, to avoid describing the target object category, we encourage the feature similarity between $p_{k}$ and $t_{k}$ to be low with $L^{T}_{prompt}$. Thus, the prompt learning loss $L_{prompt}$ is defined as below:
@@
$=$ $\mathbb{E}_{X}\left[-log(sim(u^{b}_{k},v^{b}_{k}))\right]+$
$\mathbb{E}_{X}\left[-\lambda_{T}\cdot log(1-sim(u^{b}_{k},u^{f}_{k}))\right],$
$\textit{where}\,\,u^{b}_{k}=E_{T}$ $(p_{k}),\,\,v^{b}_{k}=E_{I}(X^{b}_{k}),\,\,u^{f}_{k}=E_{T}(t_{k}).$
Here, the mask generator $S$ is fixed and $p_{k}$ is the only trainable part for loss $L_{prompt}$, and $\lambda_{T}$ is the loss weight for minimizing the similarities to the text labels. Once the above learning is complete, our prompts $p_{k}$ would represent co-occurring backgrounds for each category $k$ without requiring manually-defined background prompts, and is therefore preferable to existing CLIP-based methods [65, 42, 72, 49, 69]. In addition, our Contrastive Prompt Learning aims to capture class-associated backgrounds for segmentation purposes, rather than replacing general text templates like “a photo of {}” for classification tasks as previous prompt learning methods [83, 82, 25] do.
3.2.3 Prompt-guided Semantic Refinement
Finally, to suppress co-occurring background regions from the object mask $M$, our SemPLeS framework exploits the previously derived background prompts $p_{k}$ to perform Prompt-guided Semantic Refinement (Fig. 1 (c)). More specifically, we encourage our mask generator $S$ to produce refined masks $M^{\prime}$ by excluding the semantic knowledge embedded in the background prompts $p_{k}$, while the objectives introduced in Eq. (1) are retained to match the class labels. Hence, the refinement loss $L_{refine}$ and the total loss function $L_{total}$ are defined as follows: